OpenStudy (anonymous):

Solve: x^2-2x-24 ≤ 0 A) [-4,6] B) (-∞,-4] U [6, ∞) C) [6, ∞) D) (-∞,-4]

OpenStudy (anonymous):

x^2-2x-24 = (x-1)^2 - 25

OpenStudy (anonymous):

so simplifying you get |x-1| <= 5

OpenStudy (anonymous):

Maybe you can see now which answer is right?